OpenStudy (anonymous):

write 11 over 50 as a percent

OpenStudy (hoblos):

22%

OpenStudy (anonymous):

k

OpenStudy (anonymous):

\[\text{a as a percentage of b}=\frac{a}{b} \times \frac{100}{1}\]

OpenStudy (anonymous):

what is 1 over 10 as percent

OpenStudy (anonymous):

1/10 = 10%

OpenStudy (hoblos):

10%

OpenStudy (anonymous):

k